# The Strong Birthday Problem Revisited

### v1

@@ Abstract
We revisit the Strong Birthday Problem (SBP) introduced in DasGupta2005 . The problem is stated as follows: what is the minimum number of people we have to choose so that everyone has a shared birthday with probability at least 1/2? We derive recurrence relations to compute the probability, and further show a nice connection to the associated Stirling numbers of the second kind to derive additional recurrences. We implement the recurrences using dynamic programming as well as compute the values using the combinatorial formula, and provide numerical results.
@@ 1 Introduction
What is the probability that in a randomly chosen group of $n$ people, at least two will have the same birthday? This problem is famously known as the birthday problem (see Feller1968introductionV1 , page 33). This problem was first published by Richard von Mises in 1939 vonMises1939 , but generally attributed to Harold Davenport who came up with this about 12 years earlier, but did not publish it at the time.
@@
A simple set up for the birthday problem assumes that the birthday of any person follows a uniform distribution over the days of the year with simplifying assumptions including but not limited to no selection bias, i.e., preference for a specific birthday or month, no interference between birthdays between two persons, e.g., the hospital delays one child birth by a few hours to accommodate another emergency child birth, and no twins, etc. It can be shown that we only need 23 people to achieve a the probability more than 1/2 that at least two people share the same birthday. This is somewhat counterintuitive, given that we have 365 days in a year (disregarding the leap year), we need 366 people to attain the probability of 1, by the pigeonhole principle. A simple variant of this problem generalizes to matching birthdays of couples, that is we need husband-husband and wife-wife birthday matches. It is easy to show that (replace 365 by $365^{2}$ in the original solution to the birthday problem), we need 431 couples to achieve a probability of 1/2 or more DasGupta2005 . Yet another variant asks for the minimum number of people required so that at least two have a birthday within $d$ adjacent days with probability at least 1/2 Naus1968 . It can be shown that for $d=1,2,3,4,5$, the corresponding minimum number of people needed are $n=23,14,11,9,8$, respectively.
@@
Several variants and extensions have been proposed to the birthday problem over the past eight decades Naus1968 ; mosteller1962 ; Mathis1991 ; Nunnikhoven1992 ; DasGupta2005 ; inoue2008 . While many of them can be entertaining puzzles, their solutions, computationally efficient exact and approximation algorithms designed to compute those have been an active area of research. This stems from the fact that several real-world problems in the domains of computer science including hashing, cryptography bellare2004 , communication networks and load balancing in distributed systems eijkhout2019 , and problems in other domains including biology Green2024 and medicine Krawitz2015 , forensic science Kaye2013Beyond , statistics, social science and data science gold2024 , can be modeled as one of the variants of the birthday problem.
In this article, we consider the Strong Birthday Problem (SBP) introduced in DasGupta2005 . The problem asks to compute the minimum number of people we have to choose so that everyone has a shared birthday with probability at least 1/2. We provide both combinatorial solution (as we could not find one in the earlier publications) in Subsection 2.1, and also derive a recurrence relation to compute the probability in Subsection 2.2. Further, we show that the associated Stirling numbers of the second kind (see Subsection 2.3) has a nice connection to this problem, which we exploited to derive additional recurrence relations (see Subsection 2.4). Finally, we implement the recurrences using dynamic programming, and also computed the values using the combinatorial formula. We provide numerical results based on our implementation in Section 3, and discuss other extensions in Section 4. Throughout the paper, we also note a number of related problems that are extensions to SBP which can be considered in their right, and our solution easily generalizes to those instances.
@@ 2 The Strong Birthday Problem (SBP)
The strong birthday problem is stated as follows. In a group of $n$ people, what is the probability that each person has a shared birthday with some other person in the group? In our set up, we assume that the number of days in a year is $m$, and each person’s birthday is equally likely to be incident on one of the days of the year. That is, the probability that $i$th person’s birthday is incident on $j$th day of the year is $1/m$, for $1\leq i\leq n$ and $1\leq j\leq m$. While this is a simplifying assumption for the birthday related problems, in may be a reasonable assumption in other contexts where the problem can be recast as an instance of a variant of one of the birthday problems.
The paper DasGupta2005 gave a formula to compute the probability for SBP However, we could not find a derivation of it. Therefore, as a warm up, we give a derivation for the combinatorial formula for SBP.
@@ 2.1 A Combinatorial Formula for SBP
Let the number of people be $n$ and the number of days in a year be $m$. Let’s compute the number of ways in which $k$ singleton birthdays, i.e., a day with exactly one person’s birthday incident on it, can be formed. There are four steps in counting we need to do here.
@@
    There are $m\choose k$ ways of choosing $k$ birthdays out of $m$ days.
@@
    There are $n\choose k$ ways of choosing $k$ people out of $n$ people.
@@
    The above chosen $k$ birthdays form a bijection with the chosen $k$ people. The number of such bijections is $k!$.
@@
    Furthermore, the remaining nkn-k people can be assigned the remaining mkm-k days in $(m-k)^{n-k}$ ways.
Therefore, the number of ways of forming $k$ singleton birthdays is
$M_{m,n,k}$ $=$ ${m\choose k}{n\choose k}k!(m-k)^{n-k}.$
However, since the assignment of the remaining nkn-k people to the remaining mkm-k days can have singleton birthdays, we end up over-counting the number of singleton birthdays, which we successively correct for by using the inclusion-exclusion principle rosen2019 to finally obtain the expression $N_{m,n,k}$, for the number of ways of forming $k$ singleton birthdays:
$N_{m,n,k}$ $=$ $\sum_{j=k}^{n}(-1)^{j-k}{j\choose k}M_{m,n,j}=\sum_{j=k}^{n}(-1)^{j-k}{j\choose k}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (1)
The number of ways $n$ people can be assigned $m$ days as birthdays is $m^{n}$, which is the size of the sample space. Therefore, the probability of having $k$ singleton birthdays is
$P\!\left(m,n,k\right)$ $=$ $\frac{1}{m^{n}}N_{m,n,k}=\frac{1}{m^{n}}\sum_{j=k}^{n}(-1)^{j-k}{j\choose k}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (2)
@@
If we use $k=0$ in the above formula, it gives the probability that each person in the group of $n$ people shares her/his birthday with someone in the group:
$P\!\left(m,n,0\right)$ $=$ $\frac{1}{m^{n}}\sum_{j=0}^{n}(-1)^{j}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (3)
2.2 A Recurrence for SBP
For the strong birthday problem, we can derive a recurrence relation as follows. Let $T(j,k,n,m)$ denote the number of ways in which $n$ birthdays are distributed over $m$ possible days in a year so that there are $j$ days with 2 or more birthdays incident on them and $k$ days with singleton birthdays. Therefore, mjkm-j-k days are birthday-free. Clearly, $n\geq 2j+k$. There are three mutually exclusive cases to consider in setting up a recurrence for $T(j,k,n,m)$, as given below.
@@ 1. Case 1:
    (Add to days with 2 or more birthdays). $n$th person’s birthday is incident on one of the $j$ days with 2 or more birthdays. Then, there are T(j,k,n1,m)T(j,k,n-1,m) ways in which the first n1n-1 birthdays are distributed so that there are $j$ days with 2 or more birthdays incident on them and $k$ days with singleton birthdays. Therefore, the number of ways this can be done is jT(j,k,n1,m)jT(j,k,n-1,m).
@@ 2. Case 2:
    (Add to singleton birthdays). $n$th person’s birthday is incident on one of the $k+1$ singleton birthdays. In this case, after $n$th person is included, the number of singletons birthdays goes from $k+1$ to $k$ and the number of days with 2 or more birthday goes from j1j-1 to $j$. Therefore, the number of ways this can be done is (k+1)T(j1,k+1,n1,m)(k+1)T(j-1,k+1,n-1,m).
@@ 3. Case 3:
    (Create a Singleton birthday). $n$th person forms a singleton birthday. There are T(j,k1,n1,m)T(j,k-1,n-1,m) ways in which the first n1n-1 birthdays are distributed so that there $j$ days with 2 or more birthdays incident on them and k1k-1 days with singleton birthdays. One of the remaining mj(k1)=mjk+1m-j-(k-1)=m-j-k+1 days can be used to realize the $n$th person’s birthday. Therefore, the number of ways this can be done is (mjk+1)T(j,k1,n1,m)(m-j-k+1)T(j,k-1,n-1,m).
Hence the recurrence is given by
$T(j,k,n,m)$ $=$ $jT(j,k,n-1,m)$ (4)
$+(k+1)T(j-1,k+1,n-1,m)$
$+(m-j-k+1)T(j,k-1,n-1,m).$
@@
$T(j,k,n,m)$ $=$ $\begin{cases}0,&\text{if $j<0\lor k<0\lor n<2j+k$}\lor m<j+k;\\ 1,&\text{if $j=0\land k=0\land n=0$};\\ jT(j,k,n-1,m)\\ +(k+1)T(j-1,k+1,n-1,m)\\ +(m-j-k+1)T(j,k-1,n-1,m),&\text{otherwise.}\end{cases}$ (5)
@@
The above formula can be computed using dynamic programming. Since in the recurrence $j\in[0,n/2]$, $k\in[0,n]$ as it can grow during recursion due to the middle term in the recurrence, and $n$ ranges from 0 to $n$, the total number of states are $\operatorname{O}\left(n^{3}\right)$.Since the summation runs for $\operatorname{O}\left(n\right)$ time, the total time complexity is $\operatorname{O}\left(n^{4}\right)$ and space complexity is $\operatorname{O}\left(n^{3}\right)$. Here we assume that the complexity of addition and multiplication of large numbers is constant thereby we did not consider the bit complexity. Also, we assume that memoization and retrieval time is also constant.
2.3 Associated Stirling Numbers of the Second Kind

### v2

@@ Abstract
We revisit the Strong Birthday Problem (SBP) introduced by DasGupta DasGupta (2005), which asks for the minimum population $n$ required such that, with a probability of at least $1/2$, every individual in the group shares a birthday with at least one other person. Formally, we develop and analyze computational frameworks to determine the probability that in a group of $n$ people with birthdays distributed over $m$ days, each day either has two or more birthdays or is birthday-free. We derive both counting-based and probability-based recurrence relations to solve this problem and establish a novel connection to associated Stirling numbers of the second kind. This relationship is exploited to derive new, more efficient recurrences. Finally, we implement these recurrences using dynamic programming, provide analysis of their asymptotic complexities, and present numerical evaluations that demonstrate the practical efficiency and scalability of our proposed approaches.
@@ 1 Introduction
What is the probability that in a randomly chosen group of $n$ people, at least two will have the same birthday? This problem is famously known as the birthday problem (see Feller (1968), page 33). While first published by Richard von Mises in 1939 von Mises (1939), it is generally attributed to Harold Davenport, who conceived of the problem approximately 12 years earlier but did not publish it at the time.
@@
A standard setup for the birthday problem assumes that birthdays are uniformly distributed over the days of the year. This includes several simplifying assumptions including but not limited to no selection bias (e.g., preference for a specific birthday or month), no interference between birth events (e.g., medical delays to accommodate hospital emergencies), and no multiple births such as twins. Under these conditions, only 23 people are required to achieve a probability greater than 1/2 that at least two individuals share a birthday. This result is often considered counterintuitive given that 366 people are required to reach a probability of 1 by the pigeonhole principle (disregarding leap years). A simple variant generalizes this to matching birthdays of couples, that is, wife-wife and husband-husband birthday matches; by replacing the 365-day year with $365^{2}$ possible birthday pairs, it can be shown that 431 couples are needed to reach a probability of 1/2 or more DasGupta (2005). Another variant identifies the minimum number of people required for at least two to have birthdays within $d$ adjacent days with a probability of at least 1/2 Naus (1968). For $d=1,2,3,4,5$, the corresponding minimum numbers of people are $n=23,14,11,9,8$, respectively.
@@
Several variants and extensions have been proposed to the birthday problem over the past eight decades Naus (1968); Mosteller (1962); Mathis (1991); Nunnikhoven (1992); DasGupta (2005); Inoue and Aki (2008). While many of them can be entertaining puzzles, their solutions, computationally efficient exact and approximation algorithms designed to compute those have been an active area of research. This stems from the fact that several real-world problems in the domains of computer science including hashing, cryptography Bellare and Kohno (2004), communication networks and load balancing in distributed systems Eijkhout et al. (2019), and problems in other domains including biology Green and Mitchell (2024) and medicine Krawitz et al. (2015), forensic science Kaye (2013), statistics, social science and data science Gold and Cuellar (2024), can be modeled as one of the variants of the birthday problem.
In this article, we consider the Strong Birthday Problem (SBP) introduced in DasGupta (2005). The problem asks to compute the minimum number of people we have to choose so that everyone has a shared birthday with probability at least 1/2. While closed-form formulas offer direct solutions, in this article, we explore dynamic programming (DP) recurrences as they provide a deeper understanding of the problem’s combinatorial structure through explicit state transitions, generalize more effectively to variants with different constraints, and offer insight into the sequential evolution of probabilities as people are added. By establishing a novel connection to fundamental mathematical structures like partition theory through associated Stirling number of the second kind, we further derive efficient recurrences for SBP. Our study systematically evaluates three distinct recurrence-based approaches against the direct formula, analyzing their theoretical complexity and practical performance to provide a comprehensive guide for algorithm selection.
The remainder of this article is organized as follows. Section 2 formally defines SBP, followed by the derivation of a closed-form combinatorial formula in Section 2.1. We then present three distinct algorithmic approaches: a counting-based recurrence relation in Section 2.2, a direct probability-based recurrence in Section 2.3, and a specialized recurrence exploiting the properties of associated Stirling numbers of the second kind in Section 2.4. Section 2.5 provides a comparative analysis of the asymptotic time and space complexities for each derived method. We discuss our implementation strategy, utilizing both top-down memoization and bottom-up dynamic programming, and validate our results against the combinatorial baseline in Section 3. Finally, Section 4 explores potential extensions and identifies related problems where our proposed framework may be generalized to broader classes of occupancy problems.
@@ 2 The Strong Birthday Problem (SBP)
Problem Definition (Strong Birthday Problem): Let $m\geq 1$ be the number of days in a year, and let $n\geq 1$ be the number of people whose birthdays are independently and uniformly distributed over the $m$ days. That is, the probability that the $i$th person’s birthday is incident on the $j$th day of the year is $1/m$, for $1\leq i\leq n$ and $1\leq j\leq m$. For each day $d\in[m]$, let $b_{d}$ denote the number of people with birthdays incident on day $d$, where $\sum_{d=1}^{m}b_{d}=n$. Let $\gamma\in[0,1]$ be a probability threshold. Then the problem seeks to find the minimum number of people $n_{\min}$ such that with probability at least $\gamma$, every day has either zero birthdays or at least two birthdays, i.e., no day has exactly one birthday. Formally, we can write this as
$n_{\min}(m,\gamma)$ $=$ $\min\left\{n\geq 1:\Pr\left(\forall d\in[m]:b_{d}\neq 1\right)\geq\gamma\right\}.$ (1)
In the above formulation, setting $\gamma=1/2$, gives us the original statement of SBP.
We approach the above problem by fixing $m,n$, and then finding the value of $\gamma$, and then doing a (binary) search on $n$ to find $n_{\min}$ so that $\gamma\geq 1/2$ (or any fixed probability threshold). Therefore, for the rest of the paper, without loss of generality, we use SBP to refer to the problem of finding the probability $\gamma$ given the values of $m$ and $n$.
@@ 2.1 A Combinatorial Formula for SBP
The paper DasGupta (2005) gave a formula to compute the probability for SBP. However, we could not find a derivation of it. Therefore, as a warm-up exercise, we give a derivation of the combinatorial formula for SBP. Specifically, we show that in a population of $n$ people and a year consisting of $m$ days, the number of ways to form exactly $k$ singleton birthdays—defined as days on which exactly one birthday occurs, is given by
$N_{m,n,k}$ $=$ $\sum_{j=k}^{\min(m,n)}(-1)^{j-k}{j\choose k}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (2)
We give a proof of this in three steps: first establishing a formula for at least $j$ singletons, then applying the inclusion-exclusion principle, and finally deriving the closed form formula.
Step 1: Counting configurations with at least $j$ singleton birthdays.
Let $M_{m,n,j}$ denote the number of ways to distribute $n$ people among $m$ days such that at least $j$ specific days are singleton birthdays, i.e., each of these $j$ days has exactly one person.
To construct such a configuration:
@@
    We choose $j$ specific birthdays from $m$ total days in the year in $m\choose j$ ways.
@@
    We choose $j$ people from $n$ people whose birthdays will be incident on those selected days in $n\choose j$ ways.
@@
    We establish a bijection between the $j$ chosen people and the $j$ chosen days. The number of such bijections is $j!$.
@@
    Finally, we distribute the remaining njn-j people among the remaining mjm-j available days. This can be performed in $(m-j)^{n-j}$ ways. Note that this process may produce additional singleton days; this is consistent with our objective to account for at least $j$ singletons.
Therefore, the number of ways of forming at least $j$ singleton birthdays, i.e., the sum of configurations over all possible sets of $j$ specific days singletons, is given by
$M_{m,n,j}={m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (3)
Step 2: Applying the Inclusion-Exclusion principle.
We now derive the number of configurations containing exactly $k$ singleton birthdays, as opposed to at least $k$. Let $S$ denote the set of all possible assignments of $n$ people to $m$ days. For any subset of days $D\subseteq[m]$ with cardinality $|D|=j$, let $C_{D}$ represent the set of configurations where every day in $D$ is a singleton birthday. While $|C_{D}|$ gives the number of configurations where at least the days in the specific set $D$ are singletons, the term $M_{m,n,j}$ accounts for all possible choices of $j$ days. Specifically, $M_{m,n,j}$ represents the sum of $|C_{D}|$ over all subsets $D$ of size $j$. Therefore,
$M_{m,n,j}=\sum_{|D|=j}|C_{D}|=\binom{m}{j}\times(\text{ways to make }j\text{ specific days singletons}).$ (4)
By the inclusion-exclusion principle, the number of configurations with exactly $k$ singleton birthdays, denoted by $S_{k}$, is obtained by starting with configurations that include $k$ specific singletons, subtracting those that have $k+1$ or more singletons, and adding back over-counted terms, and so on. This systematic addition and subtraction ensures that each configuration is counted precisely once. Formally,
$|S_{k}|=\sum_{j=k}^{\min(m,n)}(-1)^{j-k}\binom{j}{k}\times(\text{configurations with at least $j$ singletons}),$ (5)
where the binomial coefficient $\binom{j}{k}$ accounts for the number of ways a configuration with $j$ singletons is overcounted among the $\binom{m}{k}$ possible sets of $k$ singletons, i.e., it counts the number of ways of choosing which $k$ of the $j$ at least singleton days are the exactly $k$ singletons. The alternating sign $(-1)^{j-k}$ performs the inclusion-exclusion adjustment. Consequently, the total number of configurations with exactly $k$ singleton birthdays, $N_{m,n,k}$, is given by
$N_{m,n,k}=\sum_{j=k}^{\min(m,n)}(-1)^{j-k}{j\choose k}M_{m,n,j}.$ (6)
Note that the upper limit of the above summation is $\min(m,n)$ because the number of singleton birthdays cannot exceed the number of available days $m$ or the total number of people $n$.
Step 3: Deriving the closed-form solution.
Substituting the expression for $M_{m,n,j}$ from Equation 2 into Equation 6, we obtain the final closed-form solution for the number of ways of forming $k$ singleton birthdays:
$N_{m,n,k}$ $=$ $\sum_{j=k}^{\min(m,n)}(-1)^{j-k}{j\choose k}M_{m,n,j}$ (7)
$=$ $\sum_{j=k}^{\min(m,n)}(-1)^{j-k}{j\choose k}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$
This completes the derivation.
We note that for $0\leq k\leq\min(n,m)$, the Equation 7 is valid, and for $k>m$ or $k>n$, $N_{m,n,k}=0$ due to the impossibility of having more singleton birthdays than people or days.
Computing the Probability: The number of ways $n$ people can be assigned $m$ days as birthdays is $m^{n}$, which is the size of the sample space. Therefore, the probability of having $k$ singleton birthdays is
$P\!\left(m,n,k\right)$ $=$ $\frac{1}{m^{n}}N_{m,n,k}=\frac{1}{m^{n}}\sum_{j=k}^{\min(m,n)}(-1)^{j-k}{j\choose k}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (8)
@@
Using $k=0$ in the above formula, we obtain the probability that each person in the group of $n$ people shares her/his birthday with someone in the group:
$P\!\left(m,n,0\right)$ $=$ $\frac{1}{m^{n}}\sum_{j=0}^{\min(m,n)}(-1)^{j}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (9)
Complexity Analysis: The formula in the Equation 9 is computed by iterating the sum from $j=0$ to $j=n$, and computing each term using the binomial coefficients, factorials and exponentials. Under the unit-cost arithmetic model, we have $O\!\left(n\right)$ iterations with $O\!\left(1\right)$ operations per iteration; therefore, the time complexity is $O\!\left(n\right)$. With arbitrary precision arithmetic, each operation has cost dependent on the bit-length of representation of the number; therefore, the time complexity is $O\!\left(n\log^{2}(mn)\right)$. The space complexity is $O\!\left(1\right)$ under unit-cost arithmetic model, and it is $O\!\left(\log^{2}(mn)\right)$ under bit-length model.
2.2 A Recurrence for SBP Based on Counting
Let $T(j,k,n,m)$ denote the number of ways in which $n$ birthdays are distributed over $m$ possible days in a year so that there are $j$ days with 2 or more birthdays incident on them and $k$ days with singleton birthdays. Therefore, mjkm-j-k days are birthday-free. Clearly, $n\geq 2j+k$. We derive the recurrence by considering the birthday of the $n$th person. Given a valid configuration with n1n-1 people, addition of $n$th person creates exactly one of the three mutually exclusive cases as given below.
@@ 1. Case 1:
    Add to days with $\geq 2$ birthdays. The $n$th person’s birthday is incident on one of the $j$ days with 2 or more birthdays. Then, there are T(j,k,n1,m)T(j,k,n-1,m) ways in which the first n1n-1 birthdays are distributed so that there are $j$ days with 2 or more birthdays incident on them and $k$ days with singleton birthdays. Therefore, the number of ways this can be done is $j\cdot T(j,k,n-1,m)$.
@@ 2. Case 2:
    Add to singleton birthdays. The $n$th person’s birthday is incident on one of the $k+1$ singleton birthdays. In this case, after $n$th person is included, the number of singletons birthdays goes from $k+1$ to $k$ and the number of days with 2 or more birthday goes from j1j-1 to $j$. Therefore, the number of ways this can be done is $(k+1)\cdot T(j-1,k+1,n-1,m)$.
@@ 3. Case 3:
    Create a singleton birthday. The $n$th person forms a singleton birthday. There are T(j,k1,n1,m)T(j,k-1,n-1,m) ways in which the first n1n-1 birthdays are distributed so that there $j$ days with 2 or more birthdays incident on them and k1k-1 days with singleton birthdays. One of the remaining mj(k1)=mjk+1m-j-(k-1)=m-j-k+1 days can be used to realize the $n$th person’s birthday. Therefore, the number of ways this can be done is $(m-j-k+1)\cdot T(j,k-1,n-1,m)$.
The above three cases partition all possible configurations giving the following recurrence:
$T(j,k,n,m)$ $=$ $j\cdot T(j,k,n-1,m)$ (10)
$+(k+1)\cdot T(j-1,k+1,n-1,m)$
$+(m-j-k+1)\cdot T(j,k-1,n-1,m).$
@@
$T(j,k,n,m)$ $=$ $\begin{cases}0,&\text{if $j<0\lor k<0\lor n<2j+k$}\lor m<j+k;\\ 1,&\text{if $j=0\land k=0\land n=0$};\\ j\cdot T(j,k,n-1,m)\\ +(k+1)\cdot T(j-1,k+1,n-1,m)\\ +(m-j-k+1)\cdot T(j,k-1,n-1,m),&\text{otherwise.}\end{cases}$ (11)
@@
Computational Complexity: The above formula can be computed using dynamic programming. Since in the recurrence $j\in[0,n/2]$, $k\in[0,n]$ as it can grow during recursion due to the middle term in the recurrence, and $n$’s value ranges from 0 to $n$, the total number of states are $O\!\left(n^{3}\right)$ where each state is computed in $O\!\left(1\right)$ time under unit-cost model and assuming that memoization and retrieval time is $O\!\left(1\right)$. Therefore, the time complexity is $O\!\left(n^{3}\right)$. Bottom-up dynamic programming using rolling array that maintains only two layers of $O\!\left(n^{2}\right)$ states, has the space complexity is $O\!\left(n^{2}\right)$ and the time complexity $O\!\left(n^{3}\right)$.
2.3 A Recurrence for SBP Based on Direct Probability Computation
In Equation 15, we count the number of configurations first and then at the end perform the division by $m^{n}$. These operations needs arbitrary precision arithmetic that adds to the bit complexity. To mitigate this, we can derive a recurrence to directly compute and track the probabilities.
Let $P(j,k,n,m)$ denote the probability that $n$ birthdays distributed over $m$ days result in $j$ days with $\geq 2$ birthdays, $k$ singleton days, and mjkm-j-k birthday-free days. We derive the recurrence by considering the birthday of the $n$th person, given the valid probability states with n1n-1 people. The probability transitions are from three different states with n1n-1 people to the state with $n$ people, as given below.
  1. Case 1:
    Add to days with $\geq 2$ birthdays. The $n$th person’s birthday is incident on one of the $j$ days with $\geq 2$ birthdays. The probability of this happening is $j/m$ which contributes $(j/m)\cdot P(j,k,n-1,m)$ to the total probability $P(j,k,n,m)$.
  2. Case 2:
    Add to singleton birthdays. The $n$th person’s birthday is incident on one of the $k+1$ singleton birthdays. In this case, after $n$th person is included, the number of singletons birthdays goes from $k+1$ to $k$ and the number of days with $\geq 2$ birthday goes from j1j-1 to $j$. The probability that of this happening is $(k+1)/m$ which contributes $(k+1)/m\cdot P(j-1,k+1,n-1,m)$ to the total probability $P(j,k,n,m)$.
  3. Case 3:
    Create a singleton birthday. The $n$th person forms a singleton birthday by choosing one of the mj(k1)=mjk+1m-j-(k-1)=m-j-k+1 available birthday-free days. The probability of this happening is (mjk+1)/m(m-j-k+1)/m which contributes $(m-j-k+1)/m\cdot P(j,k-1,n-1,m)$ to the total probability $P(j,k,n,m)$.
Summing these three mutually exclusive cases, and adding the base cases as we did in deriving Equation 11 yields the following recurrence:
$P(j,k,n,m)$ $=$ $\begin{cases}0,&\text{if $j<0\lor k<0\lor n<2j+k$}\lor m<j+k;\\ 1,&\text{if $j=0\land k=0\land n=0$};\\ \frac{j}{m}\cdot P(j,k,n-1,m)\\ +\frac{k+1}{m}\cdot P(j-1,k+1,n-1,m)\\ +\frac{m-j-k+1}{m}\cdot P(j,k-1,n-1,m),&\text{otherwise.}\end{cases}$ (16)
For SBP, the probability is computed using the following equation:
$P(m,n)$ $=$ $\sum_{j=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}P(j,0,n,m).$ (18)
We note that the two recurrences in Equation 11 and Equation 16 are related by
$P(j,k,n,m)$ $=$ $\frac{T(j,k,n,m)}{m^{n}},$ (19)
which can be proved straightforwardly by using induction (not shown here).
Computational Complexity: The bottom-up dynamic programming implementation with rolling array (only two layers of $O\!\left(n^{2}\right)$ states to store the current and previous states are needed) has the time and space complexities $O\!\left(n^{3}\right)$ and $O\!\left(n^{2}\right)$, respectively. However, since the probability recurrence uses standard floating-point arithmetic as opposed to arbitrary precision arithmetic, the arithmetic operations are done in $O\!\left(1\right)$ time which benefits from hardware-accelerated operations, giving significant speedup in practical implementations.
2.4 Associated Stirling Numbers of the Second Kind and It’s Connection to the Strong Birthday Problem
We show that associated Stirling numbers of the second kind can be used to derive novel recurrences for SBP that are both time and space efficient. Specifically, we show that the number of unique states are $O\!\left(n^{2}\right)$, and only $O\!\left(n\right)$ states need to be stored in a bottom-up rolling array implementation, giving us $O\!\left(n^{2}\right)$ time and $O\!\left(n\right)$ space comepexities. Since the recurrence is independent of $m$ during the dynamic programming computation, it enables precomputation of Stirling numbers. The recurrence we derive here has cleaner state space than the method based on counting and direct probability computation, and has better asymptotic complexity bounds.
2.4.1 Associated Stirling Numbers of the Second Kind